OpenStudy (anonymous):

Wendell deposited $2800 in a savings account. At the end of the year, he received $360 simple interest. What rate of interest did Wendell's bank pay?

OpenStudy (anonymous):

360=2800*r*1/100 r=360/28=12.8% p.a.